OpenStudy (anonymous):

What value of c solves the equation? 4power/c=1/16 c=

OpenStudy (anonymous):

@Hero

OpenStudy (alekos):

4power/c doesn't mean anything

OpenStudy (anonymous):

4 to the power of c

OpenStudy (alekos):

Is this what you mean \[4^{c} = \frac{ 1 }{ 16c }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait but without the c on the 16

OpenStudy (thread252):

I think c=1/64

OpenStudy (alekos):

OK. then what's another way of writing 1/16 ?

OpenStudy (anonymous):

I might have figured it out 4-4c=1/16-4 c=12 idk

OpenStudy (thread252):

life is so confusing

OpenStudy (alekos):

No. Just answer my question

OpenStudy (anonymous):

just 16 ?

OpenStudy (alekos):

No way. How can 1/16 = 16 ?

OpenStudy (anonymous):

sorry I switched it

OpenStudy (anonymous):

0.0625

OpenStudy (alekos):

so what's your answer?

OpenStudy (anonymous):

6.25 ?

OpenStudy (alekos):

OK. we're not really getting anywhere. Lets try this whats another way of writing 16?

OpenStudy (anonymous):

wdym

OpenStudy (alekos):

what's the square root of 16?

OpenStudy (anonymous):

4

OpenStudy (alekos):

OK, finally! So \[4^{2} = ?\]

OpenStudy (anonymous):

16

OpenStudy (alekos):

Great!! and what about \[\frac{ 1 }{ 4^{2} }\]

OpenStudy (anonymous):

0.0625

OpenStudy (alekos):

No! keep it as fraction

OpenStudy (anonymous):

ohh 1/16

OpenStudy (alekos):

OK, so now whats another way of writing \[\frac{ 1 }{ 4^{2} }\]

OpenStudy (anonymous):

1/16

OpenStudy (alekos):

We already did that. \[\frac{ 1 }{ 16 } = \frac{ 1 }{ 4^{2} } = 4^{-2}\]

OpenStudy (alekos):

so now we have \[4^{c} = 4^{-2}\] OK?

OpenStudy (anonymous):

ok thanks

OpenStudy (alekos):

we haven't finished yet. Can you see what c is equal to?